Assume that in January 2017, the average house price in a particular area was $288,600. In January 2000, the average price was $195,300.

  

What was the annual increase in selling price? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

288600=195,300*(1+r/100)^17

(288600/195,300)^(1/17)=(1+r/100)

(1+r/100)=1.0232

r=1.0232-1

=2.32%(Approx).